OpenStudy (anonymous):

A bag has 2 red marbles, 3 blue marbles and 5 white marbles. A marble is randomly selected and then placed back in the bag. You do this 8 times. What is the probability of choosing 1 red marble, 3blue marble and 4 white marbles?

OpenStudy (amorfide):

add up all the marbles to get 10 you then create fractions to represent the probability of getting each colored marble from the bag so you have 2/10 for red 3/10 for blue and 5/10 for white next you can just create a tree diagram

OpenStudy (ybarrap):

n=number of marbles (10) r=number of red marbles (2) b=number of blue marbles (3) w=number of white marbles (5) \(p_r\) = probability of drawing a red (2/10) \(p_b\) = probability of drawing a blue (3/10) \(p_w\) = probability of drawing a white (5/10) Use the multinomial distribution (see - http://en.wikipedia.org/wiki/Multinomial_distribution#Probability_mass_function ) Here is the probability for your problem $$ \large {n!\over r!b!w!}\times (p_r)^1(p_b)^3(p_w)^5 $$ Does this make sense?